the nth term of the sequence is given by tn=5n-2 if the last term is 148,how many terms are there?
Answer by sudhanshu_kmr(1152)

first term t1 = 5(1)-2 = 3
second term t2= 5(2) - 2 = 8
so, sequence is an Arithmetic Progression 3,8,13,18..........
here common difference = 5
let there are total x terms then

tx= 5x - 2
=> 148 = 5x -2
=> x = 150/5 = 30
there are 30 terms.


Sometimes there may be typing mistake in solution of a problem, so please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@gmail.com
Best of luck.......